OpenStudy (anonymous):

an estimate of the quotient of 3 3/4/ 2 4/9

OpenStudy (anonymous):

If all you want is an estimate, simply see what the mixed fractions are near. Do some quick combining of terms to make life easy. The top would be (3*4)/4 + (3/4) = 15/4 The bottom would be (2*9)/9 + (4/9) = 22/9 So the top contains something just below 4. The bottom contains about 2 1/2. So 4/2.5 is greater than 1, but less than 2. So I would estimate 1.5